OpenStudy (anonymous):

help please could i find 2 missing angle measures if i know some of the side lengths of a triangle?

OpenStudy (anonymous):

whats the question??

OpenStudy (anonymous):

that is a copy and paste of the question...i think it has something to do with cotangent, cosecant, and secant...i am suppose to make up my own example but i don't understand it that well...

jimthompson5910 (jim_thompson5910):

If you're dealing with a right triangle, then the answer is definitely yes If you're dealing with just any type of triangle (that may not be a right triangle), then the answer isn't always yes

jimthompson5910 (jim_thompson5910):

with a right triangle, you will have to know at least 2 sides to figure out the two missing angles

OpenStudy (anonymous):

sorry for taking long, computer crashed... but it is a right triangle... i am suppose to use this triangle in my example

jimthompson5910 (jim_thompson5910):

ok do you know any of the sides?

OpenStudy (anonymous):

the thing said i am suppose to make some side lengths as part of the example i make... the question that i gave is a question that a"friend" sent me(part of the lesson) and i am suppose to answer him/her by making up my own example(which includes making up the side lengths and etc.) I believe it has to deal with csc, sec, or cot but i am not familiarized with either of them.

jimthompson5910 (jim_thompson5910):

ok so you can basically write in any reasonable numbers you want?

jimthompson5910 (jim_thompson5910):

for the side lengths

OpenStudy (anonymous):

correct

jimthompson5910 (jim_thompson5910):

so what numbers did you want to use

OpenStudy (anonymous):

how about... 10 and 14?

jimthompson5910 (jim_thompson5910):

ok where do you want to put 10 and 14, go ahead and draw them in

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

if you want to know angle C, then you can say tan(C) = 10/14 tan(C) = 0.7142857 C = arctan(0.7142857) C = ?? Note: make sure you're in degree mode

jimthompson5910 (jim_thompson5910):

if you want to use cotangent, then cot(C) = 14/10 cot(C) = 1.4 1/tan(C) = 1.4 tan(C) = 1/1.4 tan(C) = 0.7142857 C = arctan(0.7142857) C = ?? but you end up getting the same thing as before

OpenStudy (anonymous):

??? confused using angle c the adjacent would be 14? and opposite is 10? 14 / 10 = 1.4 ??? how do you get 0.7142857

jimthompson5910 (jim_thompson5910):

1/tan(C) = 1.4 since cot(C) = 1.4 you then take the reciprocal of both sides so flip 1/tan(C) to get tan(C) flip 1.4 or 1.4/1 to get 1/1.4

jimthompson5910 (jim_thompson5910):

1/1.4 = 0.7142857 so 1/tan(C) = 1.4 turns into tan(C) = 0.7142857

OpenStudy (anonymous):

and that is the angle measure????

jimthompson5910 (jim_thompson5910):

not quite

jimthompson5910 (jim_thompson5910):

tan(C) = 0.7142857 C = arctan(0.7142857) C = 35.537677

jimthompson5910 (jim_thompson5910):

so angle C is roughly 35.537677 degrees

jimthompson5910 (jim_thompson5910):

now that you know that C = 35.537677, you can use this to find angle A A+C = 90 A+35.537677 = 90 A = 90 - 35.537677 A = 54.462323 So angle A is roughly 54.462323 degrees

OpenStudy (anonymous):

thank you:)

jimthompson5910 (jim_thompson5910):

sure thing